# Diverse Score Distillation

### v1

@@ 5.1 Image Generation via Score Distillation
In Fig. 4, we visualize 4 samples generated by each method across different prompts. When our optimization steps are aligned with the DDIM sampling steps (DSD), the results are equivalent to DDIM sampling. However, in the general case (DSD), where the optimization steps differ from the DDIM steps (see Appendix), slight variations do arise. Despite this, our method consistently produces generations that are more diverse and of higher fidelity compared to baseline score distillation methods.
@@ 5.2 Text to 3D Generation
We use various score distillation frameworks to obtain 4 different 3D generations from each text prompt, and visualize these in Fig. 5. Consistent with our discussion in Sec. 4, SDS [23] and ASD [20] suffer from mode-seeking behavior and produce less sharp results. While VSD can theoretically generated more results, in practice the simultaneous training of a LoRA [12] network that models the rendering distribution in VSD limits its quality.
Additionally, these methods are unstable when handling more challenging prompts. Consistent 3D is more diverse than the aforementioned methods thanks to its ODE formulation. However, the gradient direction and large CFG are harmful to the result. While SDI  [19] is capable of generating high-quality images, its diversity is limited due to the inversion-based ODE seeking.
We also quantitatively evaluate the quality and diversity of the distilled 3D shapes. For quality, we use FID  [8] and CLIP-SIM  [24]. For each prompt, FID is calculated across 100 generated samples from Stable Diffusion and the 100 rendered images from 10 3D shapes. For diversity, we propose to use LPIPS  [36] to measure the difference between different 3D shapes generated using the same prompt. For each prompt, we generate 10 3D shapes and measure the pair-wise LPIPS for the same camera. Tab. 2 demonstrates the calculated metrics, measured from 15 prompts. A smaller FID indicates that the generated 3D shapes are more similar to the 2D images produced by the DM, while a higher CLIP-SIM reflects better text-image alignment of the 3D shapes. Additionally, when the generated 3D shapes appear more distinct, the LPIPS score should be larger. As shown in Tab. 2, and consistent with the observations in Fig. 5, our method generates 3D shapes with comparable or superior quality while achieving greater diversity compared to baseline methods.
@@
SDS 263.75 27.50 0.2830
HiFA 260.10 26.28 0.3434
SDI 242.11 27.44 0.2156
Consistent 3D 256.32 26.61 0.3793
VSD 275.47 27.11 0.2849
Ours 246.25 28.62 0.3887
@@
We also investigate alternative ways of getting the $\mathbf{x}(t)$ and $\mathbf{x}(t-\delta)$ from Eq. 8: random noise (similar to ASD [20]), linear approximation (similar to Consistent3D [34] but with a corrected gradient direction), and our sampling with interpolation approximation (Sec. 3.3).
Fig. 6 and Tab. 3 show the qualitative and quantitative results, respectively. It can be observed that simulating a predefined ODE trajectory improves the generation diversity, and that our formulation leads to better results compared to a linear interpolation.
@@
Random (ASD) 261.99 27.44 0.3356
Linear 243.034 28.35 0.3643
Ours 246.25 28.62 0.3887
Table 3: Ablation. We ablate three ways of getting $\mathbf{x}(t)$ and $\mathbf{x}(t-\delta)$ (Eq. 8): random noise at each iteration (ASD [20]), linear approximation and ODE sampling approximation. Simulating the given ODE improves the overall quality and diversity.
5.3 Applying DSD to Other Tasks
Improving Single-view-to-3D Distillation. In addition to text-to-3D generation, score distillation has also been widely used for single-view-to-3D reconstruction  [18, 37] by leveraging DMs that are trained with image and camera conditions. Most methods still utilize the SDS  [23] gradient, which is prone to mode-seeking and often results in overly smooth outputs. Given an input image, the geometry of the underlying 3D shape is highly undetermined. As shown in Fig. 7, with the same DM, our method can produce diverse 3D shapes with high-frequency details.
@@ 6 Discussion
We presented DSD, a formulation for score distillation that akin to sampling from a diffusion model, can yield diverse samples via optimization, while matching/improving the fidelity compared to existing formulations. However, there are still several challenges and open questions. First, while sampling from diffusion models is (relatively) efficient, score distillation-based optimization is not and it would be interesting to explore formulations that can match the inference speed of denoising diffusion. Moreover, the photo-realism of optimized 3D representations still falls short of generations from the guiding 2D diffusion models, and it remains a challenge to bridge this gap. Finally, we primarily investigated the applications of DSD for text-to-3D and single-view reconstruction, but believe this can be more broadly applicable, for example in 3D editing [32, 7] or relighting [13].
@@ Appendix C More 2D Results
Figure 11: More 2D Distillation Results. Our method can simulate the original ODE perfectly (DSD*) when the number of DDIM steps and optimization steps are equal. Ours (DSD) also generate diverse results by stimulating the underlying ODEs, while other baselines  [19, 23, 20] have no explicit guarantee of diverse generation.

### v2

@@ 4.1 Image Generation via Score Distillation
In Fig. 4, we visualize 4 samples generated by each method across different prompts. When our optimization steps are aligned with the DDIM sampling steps (DSD), the results are equivalent to DDIM sampling. However, in the general case (DSD), where the optimization steps differ from the DDIM steps (see Appendix), slight variations do arise. Despite this, our method consistently produces generations that are more diverse and of higher fidelity compared to baseline score distillation methods. We use the official code of SDI [20] for its 2D results.
@@ 4.2 Text to 3D Generation
We use various score distillation frameworks to obtain 4 different 3D generations from each text prompt and visualize these in Fig. 5.
Consistent with our discussion in Sec. 3.4, SDS [24] and ASD [21] suffer from mode-seeking behavior and produce less sharp results. Although VSD can theoretically generate diverse results, simultaneous training of a LoRA [12] network that models the rendering distribution in VSD limits its quality and is potentially unstable. Consistent 3D is more diverse than the aforementioned methods thanks to its ODE formulation. However, the large CFG is harmful to the result. While SDI  [20] is capable of generating high-quality images, its diversity is limited due to the inversion-based ODE seeking.
We also quantitatively evaluate the quality and diversity of the distilled 3D shapes. For quality, we use FID  [8] and CLIP-SIM  [25]. For each prompt, FID is calculated across 100 generated samples from Stable Diffusion and the 100 rendered images from 10 3D shapes. For diversity, we propose to use LPIPS  [38] to measure the difference between different 3D shapes generated using the same prompt. We use generated 3D shapes per-prompt and measure the pair-wise LPIPS for the same camera. Tab. 2 demonstrates the calculated metrics, measured from 15 prompts. A smaller FID indicates that the generated 3D shapes are more similar to the 2D images produced by the DM, while a higher CLIP-SIM reflects better text-image alignment of the 3D shapes. Additionally, when the generated 3D shapes appear more distinct, the LPIPS score should be larger. As shown in Tab. 2, and consistent with the observations in Fig. 5, our method generates 3D shapes with comparable or superior quality while achieving greater diversity compared to baseline methods.
@@
SDS 270.80 27.69 0.2695
HiFA 258.99 28.29 0.3202
SDI 247.98 28.68 0.2407
Consistent 3D 270.58 28.16 0.2936
VSD 269.22 27.83 0.3336
Ours 251.40 28.97 0.4013
@@
We also investigate alternative ways of getting the $\mathbf{x}(t)$ and $\mathbf{x}(t-\delta)$ from Sec. 3.3: random noise (similar to ASD [21]), linear approximation (similar to Consistent3D [36], and our sampling with interpolation approximation (Sec. 3.3). Fig. 6 and Tab. 3 show the qualitative and quantitative results, respectively. It can be observed that simulating a predefined ODE trajectory improves the generation diversity, and that our formulation leads to better results compared to a linear interpolation.
4.3 Applying DSD to Other Tasks
Improving Single-view-to-3D Distillation. In addition to text-to-3D generation, score distillation has also been widely used for single-view-to-3D  [19, 40, 39] reconstruction by leveraging DMs that are trained with image and camera conditions. Most methods still utilize the SDS  [24] gradient, which is prone to mode-seeking and often results in overly smooth outputs. Given an input image, the geometry of the underlying 3D shape is highly undetermined. As shown in Fig. 7, with the same DM, our method can produce diverse 3D shapes with high-frequency details.
@@ 4.3 Applying DSD to Other Tasks
Random (ASD) 265.80 28.24 0.3333
Linear 274.46 28.17 0.3579
Ours 251.40 28.97 0.4013
Table 3: Ablation. We ablate three ways of getting $\mathbf{x}(t)$ and $\mathbf{x}(t-\delta)$: random noise at each iteration (ASD [21]), linear approximation and ODE sampling approximation. Simulating the given ODE improves the overall quality and diversity.
@@ 5 Discussion
We presented DSD, a formulation for score distillation that akin to sampling from a diffusion model, can yield diverse samples via optimization, while matching/improving the fidelity compared to existing formulations. However, there are still several challenges and open questions. First, while sampling from diffusion models is (relatively) efficient, score distillation-based optimization is not and it would be interesting to explore formulations that can match the inference speed of denoising diffusion. Moreover, the photo-realism of optimized 3D representations still falls short of generations from the guiding 2D diffusion models, and it remains a challenge to bridge this gap. Finally, we primarily investigated the applications of DSD for text-to-3D and single-view reconstruction, but believe this can be more broadly applicable, for example in 3D editing [34, 7] or relighting [13, 18].
@@ Appendix C More 2D Results
Figure 11: More 2D Distillation Results. Our method can simulate the original ODE perfectly (DSD*) when the number of DDIM steps and optimization steps are equal. Ours (DSD) also generates diverse results by stimulating the underlying ODEs, while other baselines  [20, 24, 21] have no explicit guarantee of diverse generation.